solve and graph y=4x+5
3y-12x=-6

Answer by mananth(16946)

-4 x + 1 y = 5
y=4x+5 .............1
12 x + 3 y = -6
y=4x-2 .............2

compare with y= mx+b
m=4 for both the lines. So they are parallel No solution